PALM-3000 Status - PowerPoint PPT Presentation

palm 3000 status n.
Download
Skip this Video
Loading SlideShow in 5 Seconds..
PALM-3000 Status PowerPoint Presentation
Download Presentation
PALM-3000 Status

play fullscreen
1 / 19
PALM-3000 Status
215 Views
Download Presentation
deiondre
Download Presentation

PALM-3000 Status

- - - - - - - - - - - - - - - - - - - - - - - - - - - E N D - - - - - - - - - - - - - - - - - - - - - - - - - - -
Presentation Transcript

  1. PALM-3000 Status A.H. Bouchez, J.E. Roberts Project 1640 Las Vegas Summit 1/28/2010

  2. PALM-3000 status The big picture: We are integrating the new PALM-3000 components in the lab, while a few remaining subsystems are in detailed design. We are on track for a pre-ship review in early July. Talk Outline: • Status by subsystem • 3388-actuator DM testing results • Integration and testing in 2010 • Commissioning plans in 2011

  3. DM3388 OAP1 TTM OAP2 DM241 FM1 Science Instrument Volume FM3 SSM2 FM2 FM4 SSM1 HOWFS & ACam Optical Bench • Detailed design of all optomechanics is complete • Currently verifying global flexure of bench with science instrument • Remaining design tasks: • Stimulus fold mirror • Enclosure • Cable routing • Cable interface boxes

  4. High-Order Deformable Mirror • Received at Caltech on Dec. 4, 2009 • Spent Dec-Jan. testing performance with a Zygo interferometer In the lab at Caltech Prior to face sheet bonding (Xinetics Inc.)

  5. HODM testing • Tested with a Zygo in I&T testbed • Plastic curtains enclose bench to allow humidity control with dessicant • Data • Unpowered and bias voltage measurements at 40-50% RH, 10-15% RH and 20-30% RH • Actuator influence functions • Actuator Linearity • Actuator Hysteresis • Actuator Creep • Reconstructor made from the influence functions was used to flatten the mirror • Other control tests include controlling to the low spatial frequency error instead of flat and maximizing the focus error to test the full mirror stroke

  6. HODM Control Result Surface Surface • Mirror in normal mode (40V bias) Temp: 75F RH: 26% • Zernikes 1-30 were removed from the bias measurement and the residual wavefront was flattened using the HODM actuators. The result is on the left. • On the right is the same measurement less Zernikes 1-30. This is what we can expect to achieve when the LODM is correcting the low-spatial-frequency errors

  7. HODM lessons • The HODM can be flattened to 10nm RMS • Simulations show that resulting wavefront will not limit contrast performance • The total stroke of the mirror is 1.06μm. • A few actuators along the bottom of the mirror have significantly lower stroke • With the LODM, the HODM will use 30% of its stroke to correct its own figure • We need to control the humidity around the HODM • Best figure between 20% and 30% RH • 40V bias rather than the nominal 50V bias gives more stroke and a better base figure at the expense of more non-linearity in the actuator gains

  8. Humidity Results • Unpowered surface figure at different humidities shows the effect of humidity on the puckers • Ideal humidity is between 13% and 30% 45% RH 74deg 13% RH 70deg 30% RH 74deg

  9. High Order Wavefront Sensor • Complete and ready for installation in lab testbed • Flexure testing ongoing. Initial results meet requirements (<0.2”/hr) • Two optics need to be replaced: Will replace and realign later in spring • Acquisition camera (not shown) needs to be aligned and tested Meets requirements (<0.05 binned pixels) See Baranec, Proc. SPIE 7015, 2008.

  10. HOWFS Detail Microlens arrays imaged with a CMOS alignment camera Shack-Hartmann spots imaged with the CCD-50

  11. Electronics Overview Legend: Fiber, high-speed switch, ethernet, AO bench cables Ethernet Pvt net Fiber Splitter User Workstation (Data Room) AO Bench Wavefront Processor (computer room) DM Racks (Cass cage pos. 2 & 3) Controller Racks (Cass cage pos. 4 & 5)

  12. DM Electronics and Cabling • Lab configuration complete and in use (serial com. to DM drivers, 99 cables, air cooled) • Current design efforts: • High-speed DM command electronics design in progress • Operational DM cable design in progress (16 cables for HODM, 1 for LODM) • Driver electronics dissipate ~6 kW. Liquid cooling design still required

  13. Wavefront Processor Computer • Wavefront Processor Computer clusters completed and deployed (JPL and Caltech) • Significant challenge has been hardware/software compatibility. Manufacturer of high-speed switch went out of business. • Have demonstrated full vector matrix multiplication using 64x64 subapertures at 1.24 kHz with current GPUs • If necessary, will upgrade to latest GPUs after pre-ship review to meet 2 kHz requirement

  14. Software Deep into coding & verifying interfaces. Expect to begin using software environment in the lab in Feb. Major components % Complete • Publish/subscribe infrastructure 100% • Command parser 95% • Includes communications interfaces • Wavefront processor 30% • Device drivers (hardware interfaces) 75% • Telemetry system 50% • Automations 25% • Graphical user interface 20%

  15. Servo Control • Flexible servo control scheme. • Detailed design of N=64 mode complete • Detailed design of other modes still in progress

  16. Palomar Infrastructure Currently defining requirements. Major components: • Power at Cassegrain focus (& AO lab) • Coolant at Cassegrain focus (& AO lab) • Telescope balance adjustments • Power in computer room • Communications between Cass, Computer Room, AO lab, and Data Room

  17. PALM-3000 Testbed Includes HODM, HOWFS, WFP computer, and 3 calibration/diagnostic sensors Will allow us to verify all electronic & software interfaces. HODM Zygo WFP computer CMOS HOWFS & ACAM HASO

  18. Lab I&T Schedule Testbed activities Dec-Jan Verify performance of HODM Feb-Mar Verify performance of HOWFS Mar-Apr Test wavefront processor, software environment May-Jun Verify high-speed wavefront control performance Jul SPIE meeting, pre-ship review

  19. PALM-3000 Schedule Overview